# Literature Review: Cyber Security Monitoring in Maritime

### v1

@@ V-B Security monitoring methods for distributed maritime systems and networks
Gyamfi et al. [35] developed an adaptive incremental passive-aggressive ML (AI-PAML) method for a network attack detection system (NADS) in an IoT-based Maritime Transportation System (MTS) environment. A resource-efficient multi-access edge computing (MEC) setting was proposed to execute the system at network edges. To avoid data saturation problems of online learning models and facilitate the updating of the detection system, the authors provided an enhanced approximate linear dependence and advanced data updating technique to remove the irrelevant data. For efficient execution, the computation-hungry algorithms of NADS were deployed through MEC servers located on the ship or onshore. Markov Chains were then applied to facilitate the execution process, avoiding congestion on a MEC server, by predicting the next available MEC server with statistical reasoning.
@@
The experimental setup for evaluating the work of Gyamfi et al. [35] was based on an actual IoT device network topology consisting of an RRH (Remote Radio Head for the IoT-MEC connection), MEC servers, and sensors-IoT device nodes in a vessel or at a port. The experimental results were obtained by using a PC running a layer 7 DoS simulator named DDOSIM. This was used to execute DDoS attacks against IoT devices with the aim being to achieve real-time responses from the developed NADS. The CICDDoS2019 dataset [^1] https://www.unb.ca/cic/datasets/ddos-2019.html was used to train the AI-PAML model. Although this dataset is quite recent and relevant for the IoT domain, it is not directly connected to the maritime domain. The authors also benchmarked the proposed model against the “Perceptron” and “Stochastic Gradient Descent” models regarding the training and prediction runtime and its latency was found to be better than the benchmark models after all models run against the DDOSIM attack.
@@
Kumar et al. [37] proposed a monitoring system that uses Cyber Threat Intelligence (CTI) extracted with the deep learning-based method, Long Short-Term Memory-based Variational Autoencoder. The system applied the Bi-directional Gated Recurrent Unit (Bi-GRU) method for detecting attacks addressing IoT devices. The main drawback of the study is the fact that no maritime-specific dataset was used for the validation of the method, but rather the TON-IoT dataset [^2] https://research.unsw.edu.au/projects/toniot-datasets which contains network traffic of consumer IoT devices. Although the study mentioned that the proposed method is computationally more expensive than traditional ML algorithms and requires specialised hardware, the computational cost assessment was not conducted in the study.
@@
Liu et al. [34] proposed a CNN-MLP based Intrusion Detection model for MTSs (Maritime Transportation Systems) which was trained via Federated Learning (FL), also called FedBatch by the authors. In the mentioned model, a Convolutional Neural Network (CNN) was utilised for data feature extractions, while Multilayer Perception (MLP) was used for attack classification to locally detect intrusion on each vessel. The study targeted IoT-based MTSs and endeavoured to address the limitations inherent in conventional learning algorithms, which consume an amount of computing resources that can not be provided in resource-constrained IoT environments. This study aims to protect data privacy by not sharing the training data obtained locally in vessels with the cloud. Instead, the models are induced locally in the remote devices, and their parameters are shared with the cloud. As the maritime environment may not enable all remote devices to submit their parameters in each learning round, the study utilises a batch federated aggregation approach. Thus, heterogeneous devices can upload their model parameters in a timely and synchronised way. The authors proposed a robust and efficient model for MTS security monitoring by training and testing their model on the NSL-KDD dataset. [^3] https://www.unb.ca/cic/datasets/nsl.html The model gained an overall performance accuracy of 88.1%, which surpassed the accuracy score of state-of-the-art CNN, i.e., 87.1%. Compared to other methods, the aforementioned method also achieved higher recall and almost the same precision on some attack types, i.e., Probe, R2L. Although much effort was devoted to researching FL-based intrusion detection in [34], one drawback of the study is the use of the NSL-KDD dataset, a general attack dataset, which lacks maritime domain specificity.
Ptasinski et. al [53] presented a Maximum Entropy Estimation based statistical ciphertext (CT) flow analysis mechanism to detect malicious flooding network attacks. The authors discussed the significant aspect of network security in the context of the Automated Digital Network System (ADNS) INC III, a critical network system of the US Navy for unclassified and secret information exchange. It interfaces between ships, shore-based Regional Network Operations and Security Centers (RNOSC) and Naval Computer and TeleCom Stations (NCTS). The authors specifically focused on protecting (Quality of Service) QoS in the CT domain against network flooding Denial of Service (DoS) attacks.
@@
The mentioned research by Ptasinski et. al [53] provided an overview of different aspects of Plaintext (PT) and CT domain, i.e., QoS, CT Pcore, community of interest (COI) PT Enclaves, and showed the feasibility of flooding attacks. The authors defined normal traffic classes within the CT Pcore and also included the network plans, i.e., application plan, PT management plan, and security plan, within their proposed architecture. Moreover, the proposed approach detected anomalies in the network by using anomaly-free packets as training data and estimated their distribution as a baseline which represented normal traffic. Then the network traffic was observed and its distribution was compared with the baseline distribution. If the observed distribution was different from the baseline one, then an anomaly was reported. The authors performed testing by simulating packet flooding attacks on the Pcore QoS on the secret enclave group of traffic classes. Furthermore, the aforementioned study provided an attack response mechanism by limiting network traffic. However, in some cases, normal traffic flow was also unnecessarily limited.
@@
In [38], Zainudin et al. introduced a decentralised trust aggregation solution which uses blockchain to provide a secure and trusted federated IDS structure. The solution offers a federated intrusion classification framework for SDN enabled marine traffic services. The scheme introduced a word wide available file system, with a network that uses blockchain for proof-of-authority (PoA). An IDS model that used FL-based collaboration was also proposed, together with the implemented, relatively simple intrusion detection and classification model. Their solution, leveraging CNN, surpassed other comparable implementations. The effectiveness of this model was evaluated using the X-IIoTID dataset, [^4] https://ieee-dataport.org/documents/x-iiotid-connectivity-and-device-agnostic-intrusion-dataset-industrial-internet-things specifically curated for IoT-related research.
@@
Tiwari et al. [42] proposed a relatively simple but still resilient model to secure sensor networks against cyber attacks and adversarial activities. Their work focuses on the maritime sector’s IoT-based systems. The framework addressed the weaknesses in the network layer of the abovementioned infrastructure. The paper introduced the data collection, pre-processing, model training, method optimisation, testing and its results. According to the results, the Enhanced Light Gradient Boosting Machine (Light-GBM) technique delivered excellent results on low computational cost and bandwidth, which is essential in the huge and complex marine IoT environment, that is also limited from the communication bandwidth’s viewpoint. The experiments and the IoT attacks classification were conducted on the Distributed Smart Space Orchestration System (DS2OS) dataset, [^5] https://www.kaggle.com/datasets/francoisxa/ds2ostraffictraces an open-source but not maritime-related dataset.
@@
Discussion. According to the analysis of 6 papers in this section, the current research on security monitoring methods for maritime networks has focused on specialised networks – IoT networks have been studied in four papers [34, 35, 37, 42], whereas two papers have targeted SDN [38] and military networks [53]. Moreover, the paper on monitoring SDN considered an IoT-specific dataset during evaluations, making IoT networks a dominant theme in relevant research. Therefore, the monitoring of other maritime network types than IoT has not received enough attention and constitutes an understudied research area.
@@
As for the network monitoring methods, supervised ML methods have been the most frequently employed (proposed in five papers [34, 37, 38, 42, 35]), with deep learning-based approaches being the most prominent of them [34, 37, 38]. On the other hand, as argued in a recent domain overview study [17], the superiority of deep learning over traditional ML methods is not yet proven in the field of network intrusion detection, and deep learning involves a significantly higher computational cost which might not always justify its use for network monitoring. These considerations are even more pronounced for maritime systems which are often resource constrained. Therefore, the research on lightweight ML methods suitable for maritime networks is another understudied area.
@@ V-C Security monitoring methods for navigational systems
Kontopoulos et al. [49] presented a distributed architecture capable of detecting spoofing and falsification attacks in streams of AIS data in real time. Average speed is measured on the shortest path between positions obtained from two consecutive AIS broadcast messages. Then, this value constitutes the basis of the spoof detection approach implemented in this study. The distributed architecture was of the Master-Workers type, with the Master node distributing the AIS messages to Worker nodes for analysis in an efficient manner incorporating load balancing techniques.
The approach by Kontopoulos et al. was validated by generating a 43,912,236 AIS messages dataset from marinetraffic.com and performing simulations by attacking the dataset in various ways. The validation dataset has not been made public. The reduction of false positives as well as the extension to more attack types was mentioned to be part of the authors’ future work.
Amro et al. [40] proposed a method for analysing navigational NMEA-0183 messages. These messages communicate the data of the various onboard sensors. The objective of the method was to identify possible anomalies and their malicious causes. For this, the authors developed relevant anomaly detection algorithms. To facilitate the analysis the authors employed a maritime cyber security testing environment that they previously developed [54] which included the NMEA-Manipulator tool enabling the generation of the anomalies by attacking sensor data. The approach included a software package for creating synthetic datasets (which were not made publicly available) including normal and attack-related NMEA messages.
@@
To develop the anomaly detection method, the authors first identified NMEA message types that affected navigation together with relevant message fields and their values. The authors also studied potential anomalies that can appear in NMEA messages and attack techniques which can trigger these anomalies, using the MITRE ATT&CK framework. [^6] https://attack.mitre.org/ Finally, the authors investigated different approaches for implementing attack and anomaly detection for NMEA messages. According to the authors, specification-based approach was best suited for that purpose, which involved the use of rules that described the normal behaviour of the system (i.e., any violation of these rules indicated an anomaly). Also, for some specific attacks that involved changing the arrival rate of NMEA messages to the system components, the authors recommended frequency-based anomaly detection that involved checking if the message rate remained within the expected boundaries. Although the study described the experiments for detecting a number of attack scenarios, precision and recall metrics were not reported in the paper (these metrics reflect the amount of false positives and false negatives).
The authors identified several limitations such as i) the possibility for false positives if the system is not well-configured, ii) the possibility of false negatives for sophisticated attacks that do not violate the rules of normal behaviour of the system, iii) not covering all the NMEA-0183 message types, iv) covering only NMEA-0183 and not other protocols, v) not considering all the possible anomalies. Finally, Amro et al. provided a discussion on possible deployment options of NMEA IDS on a vessel.
@@
Iphar, Napoli and Ray [45] analysed the weaknesses of AIS in their study. Since these weaknesses allow for falsifying and spoofing AIS messages, the authors developed a rule-based system for anomaly detection from AIS messages. In their study, the authors proposed over 900 rules or integrity items for assessing the integrity of AIS messages. The first order assessment rules considered only one field from a single AIS message, while the second order rules analysed several fields from a single message. The third order rules considered fields from several AIS messages of the same type, whereas the fourth order rules analysed several AIS messages of all types. Based on the truth values of rules and external contextual information, flags were set that were further used for calculating ship-related risks and their levels.
For evaluating the proposed method, a Python-based prototype was created which utilised a database for storing AIS data, intermediate results from computations, and contextual data. The prototype was evaluated on about 24 million real-life AIS messages collected over 6 months. In the study [45], the authors provided a detailed description of risk evaluation for 4 cases involving anomalous AIS messages (the study considered 13 cases in total). In all 4 cases, the prototype raised one or more flags. However, the study did not evaluate the system on a larger number of cases, including the assessment of a false positive rate of the system (reflected by the precision metric in ML literature). Also, the developed prototype and the dataset were not made publicly available by the authors.
@@
The study by Boudehenn et al. [15] focused on the detection of GPS spoofing. For that purpose, an anomaly detection approach was proposed which employed a One-Class Support Vector Machine (OC-SVM) and analysed NMEA messages that carried GPS information. The authors selected OC-SVM because it was easy to implement and did not require much computational resources, making it suitable for deployment on low-end hardware. During the experiments, the authors trained OC-SVM on data collected from a maritime simulator and also on real-life data collected on a small boat. For implementing GPS spoofing during the experiments, the authors utilised a device which generated a GPS signal carrying wrong GPS information.
According to the experiments, the OC-SVM based anomaly detection method featured a high recall and precision. To demonstrate the lightweight nature of the method and its low computational cost, the authors implemented the method on a Raspberry Pi based monitoring device that was suitable for deployment on real-life ships. Such evaluation of the actual computational cost of the proposed method and its suitability for real-life deployment sets this study apart from many similar studies which do not provide such information. The drawbacks of the study are the lack of publicly released experiment data and publicly available implementation of the method.
@@
Similarly to the previous paper by Boudehenn et al. [15], the study by Spravil et al. [36] focused on the detection of GPS spoofing with the analysis of relevant NMEA messages. According to the authors, the analysis of NMEA messages allowed for the development of dedicated monitoring modules which could be easily integrated into the ship’s network. To demonstrate the viability of that approach, the authors publicly released the implementation of the MANA (MAritime Nmea-based Anomaly detection) framework. The framework combined unsupervised anomaly detection methods into an ensemble, and to identify suitable methods, the authors investigated various GPS spoofing methods, with a number of methods originating from other domains than maritime (e.g., aviation and mobile phone networks). As a result of the investigation, five methods were selected for the ensemble.
To assess the performance of the MANA framework, the authors created the MARSIM dataset which contained a large number of different GPS spoofing attacks and normal GPS data. Also, the dataset was released into the public domain, [^7] https://github.com/fkie-cad/mana making the study by Spravil et al. [36] one of the very few where a public maritime-specific dataset has been utilised for maritime cyber security monitoring research. According to the experiments, the ensemble of five methods featured a higher recall than each method (the ensemble regarded the anomaly as detected if any of the five methods provided a relevant indication). As one of the limitations of their approach, the authors identified the need for well-chosen thresholds for each method to avoid false positive alerts. Another limitation of the study is the lack of performance data describing the computational cost of the MANA framework.
@@
Discussion. All five papers discussed in this section have focused on the analysis of application layer data in NMEA and AIS messages in order to detect AIS spoofing, GPS spoofing, and other navigational anomalies. As for the used methods, rule-based approaches have been suggested in three papers [49, 40, 45], whereas two papers have suggested unsupervised ML approaches [15, 36]. When employed for security monitoring, rule-based approaches are known to be very precise [17] and are thus used by many industrial security monitoring products (such as commonly used network IDS platforms like Suricata). However, creating a rule-based system requires domain experts and is a time-consuming process. One open research area is the use of rule mining algorithms in order to speed up the rule creation process.
@@
Also, the current research has focused on rule-based and unsupervised ML-based approaches, whereas papers analysed in this section have not suggested the use of supervised ML for the detection of navigational attacks and anomalies. Therefore, studying the feasibility of supervised ML for that purpose is another open research area, as is dealing with advanced issues of supervised ML like concept drift. Finally, as mentioned in [17], rule-based and ML-based security monitoring systems are complementary, with both of them having unique advantages. Therefore, building and evaluating such hybrid systems for navigational security monitoring is another future research direction. Inductive logic programming [55] is a potential hybrid approach which could be applied to automate the rule learning process in a supervised ML manner.
V-D Security monitoring methods for maritime CPSs

### v2

@@ V-B Security monitoring methods for distributed maritime systems and networks
Gyamfi et al. [40] developed an adaptive incremental passive-aggressive ML (AI-PAML) method for a network attack detection system (NADS) in an IoT-based Maritime Transportation System (MTS) environment. A resource-efficient multi-access edge computing (MEC) setting was proposed to execute the system at network edges. To avoid data saturation problems of online learning models and facilitate the updating of the detection system, the authors provided an enhanced approximate linear dependence and advanced data updating technique to remove the irrelevant data. For efficient execution, the computation-hungry algorithms of NADS were deployed through MEC servers located on the ship or onshore. The experimental results were obtained by using a PC running a layer 7 DDoS simulator named DDOSIM. This was used to execute DDoS attacks against IoT devices with the aim being to achieve real-time responses from the developed NADS. The CICDDoS2019 dataset [^1] https://www.unb.ca/cic/datasets/ddos-2019.html was used to train the AI-PAML model. Although this dataset is quite recent and relevant for the IoT domain, it is not directly connected to the maritime domain. The authors also benchmarked the proposed model against the “Perceptron” and “Stochastic Gradient Descent” models regarding the training and prediction runtime and its latency was found to be better than the benchmark models after all models run against the DDOSIM attack.
@@
Kumar et al. [42] proposed a monitoring system that uses Cyber Threat Intelligence (CTI) extracted with the DL-based method, Long Short-Term Memory (LSTM) based Variational Autoencoder. The system applied the Bi-directional Gated Recurrent Unit (Bi-GRU) method for detecting attacks addressing IoT devices. According to the authors, the proposed system featured a high performance, having a better accuracy than several other similar methods from non-maritime research literature. The main drawback of the study is the fact that no maritime-specific dataset was used for the validation of the method, but rather the TON-IoT dataset [^2] https://research.unsw.edu.au/projects/toniot-datasets which contains network traffic of consumer IoT devices. As one of the limitations, the study mentioned that the proposed DL-based method is computationally more expensive than traditional ML algorithms.
@@
Liu et al. [39] proposed a CNN-MLP based Intrusion Detection model for MTSs (Maritime Transportation Systems) which was trained via Federated Learning (FL), also called FedBatch by the authors. In the mentioned model, a Convolutional Neural Network (CNN) was utilised for data feature extractions, while Multilayer Perception (MLP) was used for attack classification to locally detect intrusion on each vessel. The study targeted IoT-based MTSs and endeavoured to address the limitations inherent in conventional learning algorithms, which consume an amount of computing resources that can not be provided in resource-constrained IoT environments. The study aimed to protect data privacy by not sharing the training data obtained locally in vessels with the cloud. Instead, the models were induced locally in the remote devices, and their parameters were shared with the cloud. The authors proposed a robust and efficient model for MTS security monitoring by training and testing their model on the generic NSL-KDD dataset. [^3] https://www.unb.ca/cic/datasets/nsl.html
@@
Ptasinski et. al [58] presented a Maximum Entropy Estimation based statistical ciphertext (CT) flow analysis mechanism to detect malicious flooding network attacks. The authors specifically focused on protecting (Quality of Service) QoS in the CT domain against DoS attacks in the context of the Automated Digital Network System (ADNS) INC III, a critical network system of the US Navy for unclassified and secret information exchange. The proposed approach detected anomalies in the network by using anomaly-free packets as training data and estimated their distribution as a baseline which represented normal traffic. For anomaly detection purposes, the network traffic was observed and its distribution was compared with the baseline distribution. Furthermore, the study provided an attack response mechanism by limiting network traffic. However, in some cases, normal traffic flow was also unnecessarily limited.
@@
Zainudin et al.[43] introduced a decentralised trust aggregation solution which used blockchain to provide a secure and trusted federated IDS infrastructure. The solution offered a federated intrusion classification framework for SDN enabled marine traffic services. The scheme utilised an interplanetary file system with a network that used blockchain for proof-of-authority (PoA). An IDS model that used FL-based collaboration was also proposed, together with the lightweight intrusion detection and classification model. The solution, leveraging CNN, surpassed other comparable implementations. The effectiveness of this model was evaluated using the X-IIoTID dataset, [^4] https://ieee-dataport.org/documents/x-iiotid-connectivity-and-device-agnostic-intrusion-dataset-industrial-internet-things specifically curated for IoT-related research but not being maritime-specific.
@@
Tiwari et al. [47] proposed a supervised ML approach to secure sensor networks against cyber attacks and adversarial activities. The work focused on maritime IoT-based systems, addressing the attack detection in the network layer of such infrastructures. The paper introduced the data collection, pre-processing, model training, method optimisation, testing and its results. According to the results, the Enhanced Light Gradient Boosting Machine (Light-GBM) technique delivered excellent results on low computational cost and network bandwidth usage, which is essential in the complex marine IoT environment, that is also limited from the communication bandwidth’s viewpoint. The experiments and the IoT attacks classification were conducted on the Distributed Smart Space Orchestration System (DS2OS) dataset, [^5] https://www.kaggle.com/datasets/francoisxa/ds2ostraffictraces an open-source but not maritime-related dataset.
Algarni et al. [35] explored the use of both supervised (LSTM-based DL) and unsupervised (Isolation Forest) ML methods for detecting network layer attacks in maritime wireless networks. The methods relied on the edge computing paradigm, i.e. processing data locally on network nodes. For comparative evaluation of supervised and unsupervised methods, the authors employed the WSN-DS dataset, [^6] https://www.kaggle.com/datasets/bassamkasasbeh1/wsnds which is not maritime specific. According to the study, LSTM-based DL outperformed Isolation Forest, demonstrating a significantly higher F1-score when detecting network attacks.
Agnew et al. [34] proposed a supervised method for monitoring naval submarine satellite communication networks. For creating the model of the satellite network, the authors utilised queuing theory. The method employed the CatBoost ML algorithm for detecting zero-day cyber attacks, training the algorithm on the data obtained from queuing analysis. For evaluation, simulations were used, and according to the authors, the method yielded a high precision, recall, and F1-score metric values. In addition, the study provided a computational cost analysis, indicating that the proposed approach is cost efficient and requires a minimal amount of CPU time.
Abbad et al. [32] suggested the use of stochastic semi-Markov processes for monitoring NMEA messages in maritime IoT networks, so that DDoS attacks in such networks could be identified in a timely fashion. The system proposed by the authors involved data collection from IoT devices and the detection of device states which reflect the likelihood of a security incident. Also, Markov chains were used to model state transitions in the network for identifying anomalies which can indicate DDoS attacks. The proposed system was evaluated by simulating normal traffic and DDoS attacks in maritime IoT systems, and according to the authors, the system demonstrated excellent performance with a low rate of false positives and false negatives.
Singh Popli et al. [31] proposed a DL-based method for monitoring underwater IoT networks which involved FL. With this distributed learning approach, a pretrained global ML model was shared with devices for further local training on local data, updating the global model with parameters received from devices. For comparative evaluation of FL with traditional centralised ML, the authors employed a non-maritime-specific CICIDS2017 dataset. [^7] https://www.unb.ca/cic/datasets/ids-2017.html However, for evaluating the attack detection capabilities of the proposed approach, the authors used a publicly available SOLIDS dataset [^8] https://drive.google.com/drive/u/0/folders/1IjFET0QOOZ4foEG-qzBLzTQKag6cUdiy developed in their lab, which is reflecting normal traffic and DDoS attacks in maritime underwater IoT networks. The study also involved the computational cost analysis for the proposed method, demonstrating its suitability for real-life networks.
@@
Discussion. According to the analysis of 10 papers in this section, the current research on security monitoring methods for maritime networks has focused on specialised networks – IoT networks have been studied in 6 papers [39, 40, 42, 47, 32, 31], whereas two papers have targeted military networks [58, 34]. Moreover, one paper on monitoring SDN networks [43] considered an IoT-specific dataset during evaluations, making IoT networks a dominant theme in relevant research. Therefore, the monitoring of other maritime network types than IoT has not received enough attention and constitutes an understudied research area.
@@
As for the network monitoring methods, supervised ML methods have been the most frequently employed (proposed in 8 papers [39, 42, 43, 47, 40, 35, 31, 34]), with DL-based approaches being the most prominent of them [39, 42, 43, 35, 31]. On the other hand, as argued in a recent domain overview study [17], the superiority of DL over traditional ML methods is not yet proven in the field of network intrusion detection, and DL involves a significantly higher computational cost which might not always justify its use for network monitoring (only one study [31] provided the relevant cost analysis). These considerations are even more pronounced for maritime systems which are often resource constrained. Therefore, the research on lightweight ML methods suitable for maritime networks is another understudied area.
@@ V-C Security monitoring methods for navigational systems
Kontopoulos et al. [54] presented a distributed architecture capable of detecting spoofing and falsification attacks in streams of AIS data in real time. Average speed is measured on the shortest path between positions obtained from two consecutive AIS broadcast messages. Then, this value constitutes the basis of the spoof detection approach implemented in this study. The architecture was of the Master-Workers type, with the Master distributing the AIS messages to Workers for analysis in an efficient manner incorporating load balancing techniques. The approach was validated by generating a 43,912,236 AIS messages dataset from marinetraffic.com and performing simulations by attacking the dataset in various ways. The validation dataset has not been made public. The reduction of false positives as well as the extension to more attack types was mentioned to be part of future work.
@@
Amro et al. [45] proposed a method for analysing navigational NMEA-0183 messages that communicate the data of the various onboard sensors. The objective was to identify possible anomalies and their malicious causes. For this, relevant anomaly detection algorithms were developed. To facilitate the analysis the authors employed a maritime cyber security testing environment that they previously developed [60] which included the NMEA-Manipulator tool enabling the generation of the anomalies by attacking sensor data. To develop the anomaly detection method, the authors first identified NMEA message types that affected navigation together with relevant message fields and their values. Potential anomalies that can appear in NMEA messages and attack techniques which can trigger these anomalies were studied using the MITRE ATT&CK framework. [^9] https://attack.mitre.org/ Finally, the authors investigated different approaches for implementing attack and anomaly detection for NMEA messages. According to the authors, specification-based approach was best suited for that purpose, which involved the use of rules that described the normal behaviour of the system (i.e., any violation of these rules indicated an anomaly). Also, for some specific attacks that involved changing the arrival rate of NMEA messages to the system components, the authors recommended frequency-based anomaly detection that involved checking if the message rate remained within the expected boundaries.
@@
Iphar et al. [50] analysed the weaknesses of AIS in their study. Since these weaknesses allow falsifying and spoofing AIS messages, a rule-based system was developed for anomaly detection from AIS messages. The authors proposed over 900 rules or integrity items for assessing the integrity of AIS messages. The first order assessment rules considered only one field from a single AIS message, while the second order rules analysed several fields from a single message. The third order rules considered fields from several AIS messages of the same type, whereas the fourth order rules analysed several AIS messages of all types. Based on the truth values of rules and external contextual information, flags were set that were further used for calculating ship-related risks and their levels. For evaluating the proposed method, a Python-based prototype was created which utilised a database for storing AIS data, intermediate results from computations, and contextual data. The prototype was evaluated on about 24 million real-life AIS messages collected over 6 months. In the study [50], the authors provided a detailed description of risk evaluation for 4 cases involving anomalous AIS messages (the study considered 13 cases in total). In all 4 cases, the prototype raised one or more flags. However, the study did not evaluate the system on a larger number of cases, including the assessment of a false positive rate of the system (reflected by the precision metric in ML literature).
@@
The study by Boudehenn et al. [15] focused on the detection of GPS spoofing. For that purpose, an anomaly detection approach was proposed which employed a One-Class Support Vector Machine (OC-SVM) and analysed NMEA messages that carried GPS information. The authors selected OC-SVM because it was easy to implement and did not require much computational resources, making it suitable for deployment on low-end hardware. During the experiments, the authors trained OC-SVM on data collected from a maritime simulator and also on real-life data collected on a small boat. For implementing GPS spoofing during the experiments, the authors utilised a device which generated a GPS signal carrying wrong GPS information. According to the experiments, the OC-SVM based anomaly detection method featured a high recall and precision. To demonstrate the lightweight nature of the method and its low computational cost, the authors implemented the method on a Raspberry Pi based monitoring device that was suitable for deployment on real-life ships. Such evaluation of the actual computational cost of the proposed method and its suitability for real-life deployment sets this study apart from many other studies.
@@
Similarly to the previous paper by Boudehenn et al. [15], the study by Spravil et al. [41] focused on the detection of GPS spoofing with the analysis of relevant NMEA messages. According to the authors, the analysis of NMEA messages allowed for the development of dedicated monitoring modules which could be easily integrated into the ship’s network. To demonstrate the viability of that approach, the authors publicly released the implementation of the MANA (MAritime Nmea-based Anomaly detection) framework. The framework combined unsupervised anomaly detection methods into an ensemble, and to identify suitable methods, the authors investigated various GPS spoofing methods, with a number of methods originating from other domains than maritime (e.g., aviation and mobile phone networks). As a result of the investigation, five methods were selected for the ensemble. To assess the performance of the MANA framework, the authors created the MARSIM dataset which contained a large number of different GPS spoofing attacks and normal GPS data. Also, the dataset was released into the public domain. [^10] https://github.com/fkie-cad/mana According to the experiments, the ensemble of five methods featured a higher recall than each method (the ensemble regarded the anomaly as detected if any of the five methods provided a relevant indication). As one of the limitations of their approach, the authors identified the need for well-chosen thresholds for each method to avoid false positive alerts.
Hossain et al. [33] proposed a semi-supervised ML approach for detecting attacks against maritime radar systems which involve manipulations of the radar image (e.g., freezing or rotating the image, adding an object to the image, etc.). One of the main purposes of semi-supervised ML is to lower the data labelling workload of human experts [61]. For the experiments, the authors used the RadarPWN dataset [^11] https://doi.org/10.5281/zenodo.6805559 which is maritime specific and publicly available (more details on the dataset can be found in [59]). For semi-supervised ML, the authors utilised the Random Forest classifier which was initially trained on a small dataset labelled by a human expert. The classifier was then used to predict the labels for unlabelled training data, and samples with the most certain predictions were added to training data. After that, the classifier was trained again on the labelled training dataset extended with newly added samples. The whole training process consisted of 10 such training iterations, extending the training dataset during each iteration with automatically labelled samples, which reduces the workload of human experts. According to the authors, the final classifier achieved a high precision, recall, and F1-score metric values.
The study by Longo et al. [37] described a number of attack scenarios for manipulating the maritime radar image which involved the misuse of the ASTERIX protocol. For detecting these attacks, the authors proposed a rule-based system which utilised candidate policies, where each candidate policy described the normal operating conditions of a radar. For making candidate policies adjustable for specific environments, the policies contained variables which were automatically set by the system. Also, if some candidate policies were not applicable in the given environment, they were automatically excluded from consideration. Selected policies were used for the real-time analysis of ASTERIX messages, and alerts were raised on policy violations. According to the study, the proposed system featured a high attack detection rate with a small number of false positives. In addition, the computational cost of the system was low, making it suitable for resource-constrained maritime environments. The experiments conducted in the study were supported by the MaCySTe testbed [38], which will be described in Section V-F.
To avoid data privacy issues of conventional ML and DL approaches that typically require centralising data on a single server, Hossain et al. [30] presented a FL-based approach with the aim to enhance the detection and classification of cyber attacks against marine radar systems. The method constructed a global model by utilising the collective learning ability of several distant clients, which involved local training at clients without sharing sensitive data. For evaluation, the RadarPWN dataset was used and partitioned into numerous subsets to represent different data sources in a collective learning environment. The authors provided a comprehensive evaluation of their approach, with high values reported for each performance metric. Additionally, the authors evaluated several DL-based approaches for the attack detection component of their system, choosing CNN as the best fit. The authors stated some limitations of their study, e.g., the need for reliable communication between clients and the central server, which might not always be practical in maritime operations due to connectivity issues.
@@
Discussion. All 8 papers discussed in this section have focused on the analysis of application layer data (e.g., NMEA or AIS messages) in order to detect AIS spoofing, GPS spoofing, image manipulation attacks against maritime radars, and navigational anomalies. As for the used methods, rule-based approaches are the most common and have been suggested in four papers [54, 45, 50, 37], whereas remaining four papers have suggested unsupervised ML [15, 41], semi-supervised ML [33], and DL [30]. When employed for security monitoring, rule-based approaches are known to be very precise [17] and are thus used by many industrial security monitoring products (such as commonly used network IDS platforms like Suricata). However, creating a rule-based system requires domain experts and is a time-consuming process. One open research area is the use of rule mining algorithms in order to speed up the rule creation process.
@@
Also, the current research has largely focused on rule-based and unsupervised ML-based approaches, whereas only two papers analysed in this section have suggested the use of supervised and semi-supervised ML for the detection of navigational attacks and anomalies. Therefore, further study of these approaches is another open research area, as is dealing with advanced ML issues like concept drift. Finally, as mentioned in [17], rule-based and ML-based security monitoring systems are complementary, with both of them having unique advantages. Therefore, building and evaluating such hybrid systems for navigational security monitoring is another future research direction. Inductive logic programming [62] is a potential hybrid approach which could be applied to automate the rule learning process in a supervised ML manner.
V-D Security monitoring methods for maritime CPS